Grade 12th passPhysical Chemistry

KHC2O4 solution is 4N when treated with NaOH. what is the normality of this solution when oxidised by KMnO4?
please explain step by step.....

Yasir Rather
7 Years agoGrade 12th pass

1 Answer

Vikas TU
7 Years ago
KHC2O4  and NaOH  is a  monobasic acid and a base respectively, thus their normality is equal their morality
 
When KHC2O4 is treated with KMnO4 redox reaction will take place
By principle of redox reaction
No of equivalent of oxidant= no of equivalent of reductant
Hint x factor of KMnO4 in acidic medium is 5 and x factor of KHC2O4 is 1
Normality=Molarity x equivalent mass
Molarity of KMnO x 5 = Molarity of K2H2C2O4 x 1
Molarity of KMnO = 4 x1 / 5 = 0.8
 
Normality = 0.8*5 = 4.